OpenStudy (anonymous):

what to is the mathematical solution to x+4x=38 and -x+6y=42?

myininaya (myininaya):

x+4x=5x=38 so x=38/5 so can you solve this equation for y: -38/5+6y=42

OpenStudy (dumbcow):

x=6 y=8

OpenStudy (anonymous):

5x=38 x=(38/5) -(38/5)+6y=42 --> 6y=16 y=16/6

OpenStudy (dumbcow):

i think he meant x+4y for the first one

myininaya (myininaya):

cow i was think in that first equation there should had been a y to

myininaya (myininaya):

but i wasnt sure

OpenStudy (anonymous):

Thank you DC you where right